# Data-efficient surrogate modeling of spectral functions using Gaussian processes: An application to the $t^{\prime}-t′′t^{\prime\prime}$-$J$ model

### 30 Mar 2026 v2

@@ Appendix B Additional Baseline Models
As an additional classical kernel-based baseline–complementary to the neural-network surrogate and our deep-kernel GP–we consider kernel ridge regression (KRR) for learning the forward map $(t^{\prime},t^{\prime\prime},J)\mapsto A(\omega)$ using the same 10% subset of the LCY training dataset, using the sci-kit learn implementation Pedregosa et al. (2011). Like GP regression, KRR learns a nonlinear map via a positive-definite kernel, but it yields a deterministic predictor obtained by solving a regularized least-squares problem in a reproducing kernel Hilbert space. In our setting, KRR serves as a simple reference point for how well a stationary kernel on the raw Hamiltonian parameters can capture the relevant structure of the spectra.
@@
Following LCY, we also include a $k$-nearest neighbors (kNN) regression baseline. kNN is a non-parametric, lazy learning method Pedregosa et al. (2011); Biau and Devroye (2015): given a training set of input–output pairs $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, the kNN prediction, $\hat{\mathbf{y}}(\mathbf{x})$, at a query point $\mathbf{x}$ is formed by a weighted average over the outputs of the $k$ closest training inputs.

### 07 May 2026 v3

@@ Appendix B Additional Baseline Models
As an additional classical kernel-based baseline–complementary to the neural-network surrogate and our deep-kernel GP–we consider kernel ridge regression (KRR) for learning the forward map $(t^{\prime},t^{\prime\prime},J)\mapsto A(\omega)$ using the same 10% subset of the LCY training dataset, using the sci-kit learn implementation (Pedregosa et al., 2011). Like GP regression, KRR learns a nonlinear map via a positive-definite kernel, but it yields a deterministic predictor obtained by solving a regularized least-squares problem in a reproducing kernel Hilbert space. In our setting, KRR serves as a simple reference point for how well a stationary kernel on the raw Hamiltonian parameters can capture the relevant structure of the spectra.
@@
Following LCY, we also include a $k$-nearest neighbors (kNN) regression baseline. kNN is a non-parametric, lazy learning method (Pedregosa et al., 2011; Biau and Devroye, 2015): given a training set of input–output pairs $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, the kNN prediction, $\hat{\mathbf{y}}(\mathbf{x})$, at a query point $\mathbf{x}$ is formed by a weighted average over the outputs of the $k$ closest training inputs.